OpenStudy (anonymous):

Use the Product and Quotient Rules as necessary to find the derivative of the given function. f (x) = 3x^3 cos(x)

OpenStudy (anonymous):

This is what I did, but I got it wrong. \[3x^3\cos(x)\] \[9x^2-\sin(x)\]

OpenStudy (anonymous):

Factor out constants: = 3 (d/dx(x^3 cos(x))) Use the product rule, d/dx(u v) = v ( du)/( dx)+u ( dv)/( dx), where u = x^3 and v = cos(x): = 3 (x^3 (d/dx(cos(x)))+cos(x) (d/dx(x^3))) The derivative of x^3 is 3 x^2: = 3 (x^3 (d/dx(cos(x)))+(3 x^2) cos(x)) The derivative of cos(x) is -sin(x): = 3 (x^3 -sin(x)+3 x^2 cos(x))

OpenStudy (anonymous):

when u take the derivative of one function, th esecond function remains as it is

OpenStudy (anonymous):

9x^2cos(x)+3x^3(-sinx)

OpenStudy (anonymous):

Corey89, that was just a big blob of idk. Can you clean that up?

OpenStudy (anonymous):

uzma, can you show the formula you used to do that?

OpenStudy (anonymous):

(f.g)'=f'g+fg'

OpenStudy (anonymous):

I'm really lost.

OpenStudy (anonymous):

where prime mean derivative

OpenStudy (anonymous):

factor out the constants, which in this case is 3, so f(x) = 3(x^3cos(x))

myininaya (myininaya):

\[(cfg)'=c(fg)'=c(f'g+fg')\]

OpenStudy (anonymous):

How did you get what f was and what g was?

myininaya (myininaya):

c=3 f=x^3 g=cos(x)

OpenStudy (anonymous):

C?!?!?!

OpenStudy (anonymous):

the product rule is (f*g) = f'*g + f*g' where ' is the derivative

myininaya (myininaya):

or you could say c=3 f=cos(x) g=x^3

OpenStudy (anonymous):

f would be 3x^3 and g would be cos x

myininaya (myininaya):

c is a constant

OpenStudy (anonymous):

I can't follow with 3 people. Can ONE person just do this step by step with nice spaces and neat and whatnot?

OpenStudy (anonymous):

:)...ok myiniaya u go

OpenStudy (anonymous):

I don't see why I need to factor out 3?

myininaya (myininaya):

you don't need to

OpenStudy (anonymous):

hammafer is really lost

OpenStudy (anonymous):

m so bad in typing :)

OpenStudy (anonymous):

Ok, so I just need to do f'g+g'f, correct?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[y=3x^3\cos(x)=cf(x)g(x) => c=3, f(x)=x^3=>g(x)=\cos(x)\] \[y'=c(f'g+fg')\]

OpenStudy (anonymous):

Do I need a constant?

myininaya (myininaya):

3 is a constant multiple

myininaya (myininaya):

you can do f=3x^3 and g=cos(x) if you want

myininaya (myininaya):

you can do f=x^3 and g=3cos(x) if you want

OpenStudy (anonymous):

I prefer doing it the second way, myininaya

myininaya (myininaya):

what 2nd way

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

you can do f=3x^3 and g=cos(x) if you want

OpenStudy (anonymous):

That one.

myininaya (myininaya):

then f'=9x^2 and g'=-sin(x) plug in to the following formula: f'g+fg'

OpenStudy (anonymous):

SO \[(9x^2)(\cos(x))+(3x^3)(-\sin(x))\]

myininaya (myininaya):

yes

OpenStudy (anonymous):

Then what?

myininaya (myininaya):

this is the derivative

OpenStudy (anonymous):

That's the answer?

myininaya (myininaya):

yes you can rearrange some stuff but it is the answer

OpenStudy (anonymous):

How do I rearrange it?

myininaya (myininaya):

\[3x^2(3\cos(x)-x \sin(x))\]

OpenStudy (anonymous):

may i say something pls?

OpenStudy (anonymous):

hamm...y dont u post ur work to be checked for error, otherwise u'll be lost more